# Structure of the free interfaces near triple junction singularities in harmonic maps and optimal partition problems

### v1

@@ 2.1. A topological lemma
Let $d\geq 3$ and $N\geq 2$. Suppose that $u=(u_{1},\dots,u_{N})\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ is a non-trivial $\gamma$-homogeneous solution of the optimal partition problem, i.e.
$u(x)=|x|^{\gamma}u\left(\frac{x}{|x|}\right)\quad\text{and}\quad\left\{\begin{aligned} -\Delta u_{i}&\leq 0,&&\text{in }\mathbb{R}^{d},\\ -\Delta(\sigma_{i}\cdot u)&\geq 0,&&\text{in }\mathbb{R}^{d},\end{aligned}\right.$
in a distributional sense, for all $i=1,\dots,N$. Suppose that all the points on the unit sphere are regular, that is
@@
is harmonic in $\mathbb{R}^{d}$.
@@
By Lemma 2.3 (applied to the sphere $\mathcal{M}=\partial B_{1}$ and $\mathcal{F}=\mathcal{F}(u)$) the value of $\sigma(x_{1})$ does not depend on the choice of the curve $\ell$. In order to prove the harmonicity of $\sigma\cdot u$ we notice that if $B_{r}(x_{0})$ is a ball such that $\mathcal{F}(u)\cap B_{r}(x_{0})=\partial\Omega_{i}\cap\partial\Omega_{j}\cap B_{r}(x_{0})$ is a smooth graph in $B_{r}(x_{0})$, then for every pair of points $x_{i}\in\Omega_{i}\cap B_{r}(x_{0})$ and $x_{j}\in\Omega_{j}\cap B_{r}(x_{0})$, we can find a curve in $\mathcal{C}(x_{i}\to x_{j})$ that crosses $\mathcal{F}(u)$ only once. Thus, by construction
$\sigma(x_{i})=-\sigma(x_{j})$
and so the function $\sigma\cdot u$ is harmonic across $\mathcal{F}(u)$. This concludes the proof of Lemma 2.4. ∎
We are now in position to prove our first main theorem.
Proof of Theorem 1.2.
In any dimension $d\geq 2$ and for any $N\geq 2$, we define:
(2.2) $A(d,N):=\inf\Big{\{}\gamma>1\ :\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)\neq\emptyset\Big{\}}.$
Since the admissible frequencies in dimension $d=2$ are classified, we have:
$A(2,N)=\frac{3}{2}\qquad\text{and}\qquad 1<A(d,N)\leq\frac{3}{2}\quad\text{when}\quad d\geq 2.$
We will show that $A(d,N)=3/2$ for every $d\geq 2$ by proving that
(2.3) $A(d-1,N)\leq A(d,N)\quad\text{for all }d\geq 2.$
In order to do this, we proceed in three steps.
Step 1. The infimum in (2.2) is reached. Indeed, suppose that there is a sequence of $\gamma_{n}$-homogeneous global solutions $u^{n}\in\mathcal{S}_{\gamma_{n}}(\mathbb{R}^{d};N)$ such that $\gamma_{n}\to A(d,N)$. Up to a rescaling, we suppose that:
$\|u^{n}\|_{L^{2}(\partial B_{1})}=1\qquad\text{and}\qquad\|\nabla u^{n}\|_{L^{2}(B_{1})}=\gamma_{n}.$
Up to extracting a subsequence, we can suppose that $u^{n}$ converges strongly in $L^{2}(B_{1};\mathbb{R}^{N})$ and $L^{2}(\partial B_{1};\mathbb{R}^{N})$, and weakly in $H^{1}(B_{1};\mathbb{R}^{N})$ to some $A(d,N)$-homogeneous global solution $u\in\mathcal{S}_{A(d,N)}(\mathbb{R}^{d},N)$.
Step 2. $\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}\neq\emptyset.$ Take any $k\in\{1,\dots,N\}$ and any connected component $\omega_{k}$ of $\Omega_{k}$. Since the component $u_{k}$ is homogeneous and non-negative in $\mathbb{R}^{d}$, its trace on $\omega_{k}\cap\partial B_{1}$ is (a multiple of) the first eigenfunction of the spherical Laplacian on $\omega_{k}\cap\partial B_{1}$ with eigenvalue
$\lambda:=\gamma(\gamma+d-2)\quad\text{where}\quad\gamma=A(d,N),$
with Dirichlet boundary conditions on $\partial\omega_{k}\cap\partial B_{1}$. In particular, by the Faber-Krahn inequality, the area of each $\omega_{k}\cap\partial B_{1}$ is bounded from below by a constant depending only on $\gamma$ and $d$, so the number of connected components of $\Omega_{k}$ is finite. Let $\widetilde{N}$ be the number of all connected components of $\mathbb{R}^{d}\setminus\mathcal{F}(u)$. We define the $\gamma$-homogeneous $\widetilde{u}\in\mathcal{A}(\mathbb{R}^{d};\widetilde{N})$ as the vector whose components are the restrictions of $u$ to the connected components of $\Omega_{k}$, $k=1,\dots,N$. It is immediate to check that $\widetilde{u}$ is still a global solution of the optimal partition problem, i.e. $\widetilde{u}\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};\widetilde{N})$. We now claim that $\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}\neq\emptyset$. If not, since by construction all $\widetilde{\Omega}_{j}:=\{\widetilde{u}_{j}>0\}$, $1\leq j\leq\widetilde{N}$ are connected, we can apply Lemma 2.4; thus, there is a map $\sigma:\{1,\dots,\widetilde{N}\}\to\{-1,1\}$ such that $\sigma\cdot\widetilde{u}$ is harmonic and $A(d,N)$-homogeneous in $\mathbb{R}^{d}$. Thus $A(d,N)$ is an integer and, since $A(d,N)>1$, we get that $A(d,N)\geq 2$. This is a contradiction with the bound $A(d,N)\leq 3/2$.
Step 3. Dimension reduction. From the previous step we know that there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$. Without loss of generality, we can assume that $x_{0}=\vec{\bm{e}}_{1}$. By the definition of $A(d,N)$ we know that
$A(d,N)\leq\gamma(u,\vec{\bm{e}}_{1}).$
On the other hand, the homogeneity of $u$ implies that
$\gamma(u,t\vec{\bm{e}}_{1})=\gamma(u,\vec{\bm{e}}_{1})\quad\text{for every}\quad t>0.$
In particular, by the upper semicontinuity of $\gamma(u,\cdot)$ and the homogeneity of $u$, we have
$A(d,N)=\gamma(u,0)\geq\lim_{t\to 0^{+}}\gamma(u,t\vec{\bm{e}}_{1})=\gamma(u,\vec{\bm{e}}_{1}).$
Thus, we get
$A(d,N)=\gamma(u,0)=\gamma(u,\vec{\bm{e}}_{1}),$
but this implies that $u$ is invariant in the direction of $\vec{\bm{e}}_{1}$, that is:
$u(x+t\vec{\bm{e}}_{1})=u(x)\quad\text{for every}\quad t\in\mathbb{R}.$
But then the nontrivial function
$\widetilde{u}(x_{2},\dots,x_{d}):=u(0,x_{2},\dots,x_{d}),$
belongs to $\mathcal{S}_{A(d;N)}(\mathbb{R}^{d-1};N)$. This implies that $A(d-1;N)\leq A(d;N)$, thus proving (2.3) and so (2.2).
In order to conclude the proof of Theorem 1.2, we are left to classify the $\nicefrac{{3}}{{2}}$-homogeneous global solutions. Let $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$. By Step 3 there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$. But then, reasoning as in Step 3, we have that
$\gamma(u,0)=\gamma(u,x_{0})=\frac{3}{2}$
@@
and that $u$ is invariant in the direction of $x_{0}$. Up to a rotation of the coordinate system, we can suppose that $x_{0}=\vec{\bm{e}}_{1}$. Then $\widetilde{u}(x_{2},\dots,x_{d}):=u(0,x_{2},\dots,x_{d})$ is a $\nicefrac{{3}}{{2}}$-homogeneous global solution in $\mathbb{R}^{d-1}$, i.e. it belongs to $\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d-1};N)$. By repeating this argument in every dimension $d-k$, $k\geq 1$, we finally get that, up to a rotation of the coordinate axes, we have
@@
3. Optimal partition of the sphere
The purpose of this section is to prove Theorem 1.5, essentially basing ourselves on the results obtained in Section 2, namely Theorem 1.2 (and in [CTV05b, HHT09]). By [CTV05b, Theorem 3.1] we know that there exists $\phi\in H^{1}(\partial B_{1};\Sigma_{3})\cap C^{0,1}(B_{1};\mathbb{R}^{3})$ such that $\phi_{i}\not\equiv 0$ and $(\omega_{1},\omega_{2},\omega_{3})$, defined as $\omega_{i}:=\{\phi_{i}>0\}$, achieve
(3.1) $\mathcal{L}_{3}:=\min\left\{\max_{i=1,2,3}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\}.$
Moreover, [CTV05b, Theorem 3.1] it is also proved that $\phi\in\mathcal{S}(\partial B_{1},3)$, in the sense that
$-\Delta_{\partial B_{1}}\phi_{i}&\leq 0,&&\text{in }\partial B_{1},\\ -\Delta_{\partial B_{1}}(\sigma_{i}\cdot\phi)&\geq 0,&&\text{in }\partial B_{1}.\end{aligned}\right.$
Now we recall [HHT09, Remark 3.14] which says that all the first eigenvalues are equal, namely $\lambda_{1}(\omega_{i})=\mathcal{L}_{3}$ for all $i=1,2,3$; we then let $\gamma>0$ be such that
$\gamma(\gamma+d-2)=\mathcal{L}_{3}.$
We now consider the $\gamma$-homogeneous extensions of $\phi_{i}$, that is $\varphi_{i}(r,\theta):=r^{\gamma}\phi_{i}(\theta)$. One can easily check that $\varphi\in H^{1}_{\textup{loc}}(\mathbb{R}^{d};\Sigma_{3})\cap C^{0,1}_{\textup{loc}}(\mathbb{R}^{d};\mathbb{R}^{3})$ belongs to the class $\mathcal{S}_{\gamma}(\mathbb{R}^{d},3)$, i.e.
$-\Delta\varphi_{i}&\leq 0,&&\text{in }\mathbb{R}^{d},\\ -\Delta(\sigma_{i}\cdot\varphi)&\geq 0,&&\text{in }\mathbb{R}^{d}.\end{aligned}\right.$
At this point, the proof of Theorem 1.5 follows easily. Indeed, one can test (3.1) with the $Y$-configuration and directly obtain that
$\mathcal{L}_{3}\leq\frac{3}{2}\left(\frac{3}{2}+d-2\right),$
thus implying that $\gamma\leq\nicefrac{{3}}{{2}}$. On the other hand, since $N=3$, by applying Theorem 1.2 to $\varphi$, we obtain that $\gamma=\nicefrac{{3}}{{2}}$ and that $\varphi$ coincides with the $Y$-configuration.
@@ 6. Regularity of the free boundary around points of frequency $\nicefrac{{3}}{{2}}$
In this section we complete the proof of Theorem 1.1 by proving the regularity of the free interface $\mathcal{F}(u)$ around points of frequency $\nicefrac{{3}}{{2}}$. The main ingredients are the classification of the $3/2$-homogeneous blow-ups (Theorem 1.2) and the epiperimetric inequality Theorem 1.3.
@@ 6.1. Rate of convergence of the blow-up sequence
We will show that we can apply the epiperimetric inequality to the rescalings of $u$ of the form $u^{r,x_{0}}(x)=r^{-\nicefrac{{3}}{{2}}}u(x_{0}+rx)$. In order to do so, we need to show that the conditions of Theorem 1.3 are fulfilled uniformly, at all points $x_{0}$ and at all small scales $r$, by multiples of $u^{r,x_{0}}$. We start with the following
@@
Let $u\in\mathcal{M}(D;N)$ be such that $\gamma(u,0)=\nicefrac{{3}}{{2}}$. Then, for all $\varepsilon>0$ there exists $\rho,r_{0}>0$ (depending on $d$, $N$ and $\varepsilon$) such that for all $z\in B_{\rho}\cap\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and for all $r<r_{0}$ there exists $Y\in\mathcal{M}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$ such that
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u^{z,r},Y)\right\lVert_{L^{\infty}(B_{\nicefrac{{1}}{{2}}})}\leq\varepsilon,$
where $u^{z,r}(x):=H(u,z,r)^{-\nicefrac{{1}}{{2}}}u(rx+z)$.
@@
(6.4) $\sum_{i=1}^{N}\int_{B_{1}}|\nabla u^{z_{n},r_{n}}_{i}|^{2}\,\mathrm{d}x=\mathcal{N}(u^{z_{n},r_{n}},0,1).$
@@
(6.5) $u^{z_{n},r_{n}}\to u_{0}\quad\text{strongly in }H^{1}(B_{1};\Sigma_{N})~{}\text{and uniformly in }B_{\nicefrac{{1}}{{2}}},$
for some $u_{0}\in H^{1}(B_{1};\Sigma_{N})$. Moreover, since $u^{z_{n},r_{n}}\in\mathcal{M}(B_{1};N)$ for all $n$, also $u_{0}\in\mathcal{M}(B_{1};N)$. We now observe that, since $z_{n}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and by monotonicity, we have
$\frac{3}{2}=\mathcal{N}(u,z_{n},0^{+})\leq\mathcal{N}(u,z_{n},Rr_{n})=\mathcal{N}(u^{z_{n},r_{n}},0,R),$
@@
By (6.5) and scaling properties, we have that
$\mathcal{N}(u_{0},0,R)=\lim_{n\to\infty}\mathcal{N}(u^{z_{n},r_{n}},0,R)=\lim_{n\to\infty}\mathcal{N}(u,z_{n},Rr_{n}),$
for all $R>0$. Therefore, by upper semicontinuity and monotonicity of the frequency function and Lemma 2.4, we have that for all $R,\tau>0$ sufficiently small, there holds
@@
Let $\alpha:=\varepsilon(d+1)$, with $\varepsilon>0$ being as in Theorem 1.3. Let $D\subseteq\mathbb{R}^{d}$ be open and let $u\in\mathcal{M}(D;N)$. Then, for all compact $K\subseteq D$ there exists $R>0$ such that for all $x_{0}\in K$ satisfying $\gamma(u,x_{0})=\nicefrac{{3}}{{2}}$, there exists $C_{\textup{rate}}>0$ depending on $d$ and $\mathcal{N}(u,x_{0},R)$ such that
@@
As a corollary, we also obtain the following
@@
Let $\bar{\alpha}:=2\alpha/(\alpha+3)$, where $\alpha$ is as in Proposition 6.3 and let $u\in\mathcal{M}(D;N)$. Then, for all compact $K\subseteq D$ there exists $r_{0}>0$ (depending on $d$ and $K$) and $C_{\textup{osc}}>0$ (depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$) such that
@@
for $r$ sufficiently small depending on $d$ and $\mathrm{d}_{\mathcal{H}}(K,\partial D)$ and $C>0$ depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$. On the other hand, if $L_{K}>0$ denotes the Lipschitz constant of $u$ in $K$, we have that

### v2

@@ 3.2. Classification of the blow-ups of frequency $3/2$
Let $d\geq 3$ and $N\geq 2$. Suppose that $u=(u_{1},\dots,u_{N})\in\mathcal{S}_{\gamma}(\mathbb{R}^{d};N)$ is a non-trivial $\gamma$-homogeneous function. Suppose that all the free boundary points on the unit sphere are regular, that is
@@
is harmonic in $\mathbb{R}^{d}$. In particular, $\gamma$ is an integer.
@@
By Lemma 3.3 (applied to the sphere $\mathcal{M}=\partial B_{1}$ and $\mathcal{F}=\mathcal{F}(u)$) the value of $\sigma(x_{1})$ does not depend on the choice of the curve $\ell$. In order to prove the harmonicity of $\sigma\cdot u$ we notice that if $B_{r}(x_{0})$ is a ball such that $\mathcal{F}(u)\cap B_{r}(x_{0})=\partial\Omega_{i}\cap\partial\Omega_{j}\cap B_{r}(x_{0})$ is a smooth graph in $B_{r}(x_{0})$, then for every pair of points $x_{i}\in\Omega_{i}\cap B_{r}(x_{0})$ and $x_{j}\in\Omega_{j}\cap B_{r}(x_{0})$, we can find a curve in $\mathcal{C}(x_{i}\to x_{j})$ that crosses $\mathcal{F}(u)$ only once. Thus, by construction $\sigma(x_{i})=-\sigma(x_{j})$ and so the function $\sigma\cdot u$ is harmonic across $\mathcal{F}(u)$. This concludes the proof of Lemma 3.4. ∎
Proposition 3.5.
Let $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$. Then, up to relabeling the components of $u$, we have that $u_{j}\equiv 0$ for $j\geq 4$, and, up to a rotation of the coordinate system, $(u_{1},u_{2},u_{3})$ is of the form $(u_{1},u_{2},u_{3})=cY$ for some constant $c\in\mathbb{R}$ and $Y=(Y_{1},Y_{2},Y_{3})$ as in (1.3).
Proof.
Let $u\in\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};N)$. By Lemma 3.4, there is a point $x_{0}\in\left(\mathcal{F}(u)\setminus\mathcal{F}_{1}(u)\right)\cap\partial B_{1}$. But then, the upper semicontinuity of $\gamma(u,\cdot)$ and the homogeneity of $u$ give
$\frac{3}{2}=\gamma(u,0)\geq\lim_{t\to+\infty}\gamma(u,tx_{0})=\gamma(u,x_{0}).$
Since, by [ST15], $\nicefrac{{3}}{{2}}$ is the minimal frequency, we get that
$\frac{3}{2}=\gamma(u,0)=\gamma(u,x_{0}),$
@@
which implies that $u$ is invariant in the direction of $x_{0}$. Up to a rotation of the coordinate system, we can suppose that $x_{0}=\vec{\bm{e}}_{1}$. Then $\widetilde{u}(x_{2},\dots,x_{d}):=u(0,x_{2},\dots,x_{d})$ is a $\nicefrac{{3}}{{2}}$-homogeneous global solution in $\mathbb{R}^{d-1}$, i.e. it belongs to $\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d-1};N)$. By repeating this argument in every dimension $d-k$, $k\geq 1$, we finally get that, up to a rotation of the coordinate axes, we have
@@
We notice that as a consequence of Proposition 3.5, we obtain the uniqueness of the minimizer (in any dimension) to the optimal 3-partition of the sphere for the min-max problem studied in [CTV05a, HHT09, HHT10, ST15].
Corollary 3.6.
The $Y$-configuration is the unique minimizer of
$\mathcal{L}_{3}:=\min\left\{\max_{i\in\{1,2,3\}}\lambda_{1}(\omega_{i})\colon\omega_{i}\subseteq\partial B_{1}~{}\text{open, connected and s.t. }\omega_{i}\cap\omega_{j}=\emptyset~{}\text{for }i\neq j\right\},$
where $\lambda_{1}(\omega_{i})$ is the first eigenvalue of the spherical Dirichlet Laplacian on $\omega_{i}$.
Proof.
Let $(\omega_{1},\omega_{2},\omega_{3})$ be a triple achieving the minimum (see for instance [HHT09]). We know that $\omega_{i}\subset\partial B_{1}$ is measurable and $\mathcal{H}^{d-1}(\omega_{i})>0$, for every $i$. Moreover, if $\phi_{i}$ is a first normalized eigenfunction on $\omega_{i}$, then by [HHT09, Theorem 3.4], there are constants $a_{i}\geq 0$, $i=1,2,3$, such that $\psi:=(a_{1}\phi_{1},a_{2}\phi_{2},a_{3}\phi_{3})$ satisfies
$-\Delta\psi_{i}&\leq\mathcal{L}_{3}\psi_{i},&&\text{in }\partial B_{1},\\ -\Delta\Big{(}\psi_{i}-\sum_{j\neq i}\psi_{j}\Big{)}&\geq\mathcal{L}_{3}\Big{(}\psi_{i}-\sum_{j\neq i}\psi_{j}\Big{)},&&\text{in }\partial B_{1}.\end{aligned}\right.$
By [HHT09, Theorem 3.4] that $(a_{1},a_{2},a_{3})\neq(0,0,0)$. We now claim that $a_{i}>0$ for all $i$. Suppose by contradiction that $a_{3}=0$, so that also $\psi_{3}=0$. By, the equation above, we get that $-\Delta(\psi_{1}-\psi_{2})=\mathcal{L}_{3}\,(\psi_{1}-\psi_{2})$ on $\partial B_{1}$. By the unique continuation principle for eigenfunctions, we get that $\mathcal{H}^{d-1}(\{\psi_{1}-\psi_{2}=0\})=0$, which leads to a contradiction since $\mathcal{H}^{d-1}(\omega_{3})>0$. Now, in view of [HHT09, Remark 3.14 (d)] and [ST15, Theorem 1.10], we know that $\mathcal{L}_{3}=\lambda_{1}(\omega_{i})=\frac{3}{2}\left(\frac{3}{2}+d-2\right)$ for all $i$. Thus, the $\nicefrac{{3}}{{2}}$ homogeneous extension $u$ of $\psi$ belongs to the class $\mathcal{S}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$, so the conclusion follows from Proposition 3.5. ∎
@@ 6. Regularity of the free boundary around points of frequency $\nicefrac{{3}}{{2}}$
In this section we complete the proof of Theorem 1.1 by proving the regularity of the free interface $\mathcal{F}(u)$ around points of frequency $\nicefrac{{3}}{{2}}$. The main ingredients are the classification of the $\nicefrac{{3}}{{2}}$-homogeneous blow-ups (see Proposition 3.5) and the epiperimetric inequality Theorem 1.2.
@@ 6.1. Rate of convergence of the blow-up sequence
We will show that we can apply the epiperimetric inequality to the rescalings of $u$ of the form $u^{r,x_{0}}(x)=r^{-\nicefrac{{3}}{{2}}}u(x_{0}+rx)$. In order to do so, we need to show that the conditions of Theorem 1.2 are fulfilled uniformly, at all points $x_{0}$ and at all small scales $r$, by multiples of $u^{r,x_{0}}$. We start with the following.
@@
Let $u\in\mathcal{M}(B_{1};N)$ be such that $\gamma(u,0)=\nicefrac{{3}}{{2}}$. Then, for all $\varepsilon>0$ there exists $\rho,r_{0}>0$ (depending on $d$, $N$ and $\varepsilon$) such that for all $z\in B_{\rho}\cap\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and for all $r<r_{0}$ there exists $Y\in\mathcal{M}_{\nicefrac{{3}}{{2}}}(\mathbb{R}^{d};3)$ such that
$\left\lVert\mathrm{d}_{\Sigma_{N}}(u_{z,r},Y)\right\lVert_{L^{\infty}(B_{\nicefrac{{1}}{{2}}})}\leq\varepsilon,$
where $u_{z,r}(x):=H(u,z,r)^{-\nicefrac{{1}}{{2}}}u(rx+z)$.
@@
(6.4) $\sum_{i=1}^{N}\int_{B_{1}}|\nabla(u_{z_{n},r_{n}})_{i}|^{2}\,\mathrm{d}x=\mathcal{N}(u_{z_{n},r_{n}},0,1).$
@@
(6.5) $u_{z_{n},r_{n}}\to u_{0}\quad\text{strongly in }H^{1}(B_{1};\Sigma_{N})~{}\text{and uniformly in }B_{\nicefrac{{1}}{{2}}},$
for some $u_{0}\in H^{1}(B_{1};\Sigma_{N})$, $u_{0}\not\equiv 0$. Moreover, since $u_{z_{n},r_{n}}\in\mathcal{M}(B_{1};N)$ for all $n$, also $u_{0}\in\mathcal{M}(B_{1};N)$. We now observe that, since $z_{n}\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and by monotonicity, we have
$\frac{3}{2}=\mathcal{N}(u,z_{n},0^{+})\leq\mathcal{N}(u,z_{n},Rr_{n})=\mathcal{N}(u_{z_{n},r_{n}},0,R),$
@@
By the continuity of $x\mapsto\mathcal{N}(u,x,r)$, the monotonicity of $r\mapsto\mathcal{N}(u,x,r)$ and Lemma 3.4, we have that for all $R,\tau>0$ sufficiently small, there holds
@@
Let $\alpha:=\varepsilon(d+1)$, with $\varepsilon>0$ being as in Theorem 1.2 and let $u\in\mathcal{M}(B_{1};N)$. Then, for all compact $K\subseteq B_{1}$ there exists $R>0$ such that for all $x_{0}\in K\cap\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, there exists $C_{\textup{rate}}>0$ depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,\mathrm{d}_{\mathcal{H}}(K,\partial B_{1}))$ such that
@@
As a corollary, we also obtain the following.
@@
Let $\bar{\alpha}:=2\alpha/(\alpha+3)$, where $\alpha$ is as in Proposition 6.3 and let $u\in\mathcal{M}(B_{1};N)$. Then, for all compact $K\subseteq B_{1}$ there exists $r_{0}>0$ (depending on $d$ and $K$) and $C_{\textup{osc}}>0$ (depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,\mathrm{d}_{\mathcal{H}}(K,\partial B_{1}))$) such that
@@
for $r$ sufficiently small depending on $d$ and $\mathrm{d}_{\mathcal{H}}(K,\partial B_{1})$ and $C>0$ depending on $d$ and $\sup_{x\in K}\mathcal{N}(u,x,r_{0})$. On the other hand, if $L_{K}>0$ denotes the Lipschitz constant of $u$ in $K$, we have that